Circular No. 8913

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

HDE 269858 = RADCLIFFE 127 = SANDULEAK −69^{o}220 N. R. Walborn, Space Telescope Science Institute (STScI), commu-

nicates that spectroscopy during 2008 Jan. 7–13 by R. C. Gamen at the CASLEO 2.15-m telescope in Argentina, and by N. I. Morrell at the Las Campanas Observatory 2.5-m du Pont reflector in Chile, reveals that R127, which was discovered to have entered a classical “luminous blue variable” (LBV) outburst from its Ofpe/WN9-spectral-type quiescent state in 1982 (IAUC 3767; Stahl et al. 1983, A.Ap. 127, 49), is now returning blueward from the S-Dor peculiar A-supergiant maximum state (Stahl and Wolf 1986, A.Ap. 154, 243; Wolf et al. 1988, A.Ap. Suppl. 74, 239). O. Stahl (University of Heidelberg) provides prior data showing that the spectrum had already begun to change in 1999–2002, coinciding with a sharp 1.5-mag drop in the AAVSO visual lightcurve extracted by T. Szeifert (European Southern Observatory). The AAVSO data show the object at mag ∼ 10 through the end of 2007, compared to ∼ 8.5 at maximum (early to mid-1990s) and perhaps 11.5 at the original (pre-1982) minimum. However, the V magnitude on 2008 Jan. 23–24, determined by Gamen from a “PROMPT” Cerro Tololo observation, is 10.7, fainter than at any time since 1982, indicating that a rapid decline is now occurring. The 2008 spectra display marked He I P-Cyg profiles with Si III and N II lines, but much weakened Fe II, corresponding to the intermediate peculiar B-supergiant state. Also, the blue Balmer lines have developed a second, broader (FWHM ∼ 90 km/s, possibly itself double) and more highly blueshifted (−235 km/s relative to the emission peak) absorption component, as compared to the maximum component (FWHM = 25 km/s, velocity −80 km/s) still visible. Next, R127 may return to the Ofpe/WN9 state, with He II and N III features (Walborn 1977, Ap.J. 215, 53; Walborn 1982, Ap.J. 256, 452), or it may develop dust and an infrared “catastrophe” (in which all the shorter-wavelength flux shifts to the infrared). Further spectroscopic and photometric observations are essential during the remainder of the current and subsequent Large Magellanic Cloud observing seasons, to take full advantage of this unique opportunity to follow a decades-long LBV outburst in detail. (Additional current members of this investigation team are H. E. Bond, STScI; I. D. Howarth, University College, London; and R. M. Humphreys and K. Davidson, University of Minnesota).

Corrigendum. On IAUC 3767, line 4, for hydrogen-rich read nitrogenrich

2008 January 25  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2008 CBAT  Daniel W. E. Green